In the coordinate plane the point X(4,1)is translated to the point X(5,-1) under the same translation the points Y(2,5)and Z(0,3) are translated to Y and Z. What are the coordinates of Y and Z


1
Expert's answer
2021-09-22T00:46:12-0400

The translation vector is:

"T = X2 - X1 = (5-4, - 1-1) = (1, - 2).\\"

To obtain now points one should add this vector to the initial points:

"Y2 = Y1 + T = (2 + 1, 5 - 2) = (3,3)\\
Z2 = Z1 + T = (0+1, 3-2) = (1, 1)\\"

Answer:

"(3,3), (1,1)"